OpenStudy (anonymous):

Find the derivative of the function. f(t) = t^(3/2)log9(sqrt(t + 5)) the 9 is a small number after the log

OpenStudy (anonymous):

\[f(t)=t^{3/2}\log_{9}(\sqrt{t+5}) \]

OpenStudy (anonymous):

product rule

OpenStudy (anonymous):

okiay could you help me with that

OpenStudy (anonymous):

the product rule i get but just with this problem im confused

OpenStudy (anonymous):

\[f(t)=\frac{1}{2}t^{3/2}\log_{9}(t+5)\]

OpenStudy (anonymous):

lets say y=log(base 9)(t+5) y=ln(t+5)/ln(9) change of base formula \[y'=\frac{1}{\ln(9)(t+5)}\]

OpenStudy (anonymous):

f(t)=(1/2)t^(3/2)y f'(t)=(1/2)((3/2)t^(1/2)y+y't^(3/2)

OpenStudy (anonymous):

now we just substitue the values back in

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f'(t)=\frac{1}{2}(\frac{3}{2}t^{\frac{1}{2}}\log_{9}(t+5)+\frac{t^{\frac{3}{2}}}{\ln(9)(t+5)})\]

OpenStudy (anonymous):

the person that gave u this problem is a jerk